Another rough day for our little Jack is coming to a close. As I have mentioned he has struggled for the last couple of days maintaining good blood pressures and getting rid of fluid. The doctors were unable to find anything conclusive in the cath lab last night so in an effort to exhaust all resources they did a CT scan this morning. The surgeon was at a loss as to what was causing the hypotension and all of the tests they performed were inconclusive. In an attempt to find out the source of the issue, Dr. Hawkins operated on Jack this evening. He started about 5:00 p.m. and met with us around 7:00 p.m. He told us that his heart is pumping very well, that the corrections made during the initial surgery are functioning as expected, and that by all measures he should not be having the problems that he is experiencing. In the initial operation he used a substance called "gel foam" to stop the bleeding around the vessels. Sometimes, because of the amount that is used, the foam puts strain on the heart and doesn't allow it to be as efficient as possible. Dr. Hawkins suspected that this may be the culprit, however, after removal of the "gel foam" he expected to see a fairly rapid increase in the blood pressure which unfortunately didn't occur. The team of doctors remain at a loss as to what is behind the lack of progress. We continue to be hopeful, and are careful to report that in the hours following tonights operation, Jack has better blood pressure and seems to be responding to the care given him. Once again, our hearts are full of thanks to all of you who continue to keep Baby Jack in your prayers.